OpenStudy (anonymous):

sec ^ 6 X - csc ^ 6 X

OpenStudy (anonymous):

what u hav to find in this ? dear

OpenStudy (anonymous):

I'm just an elementary child, I don't know how to solve this, can you show me the formula??

OpenStudy (anonymous):

Factor that expression

OpenStudy (anonymous):

got it dear.....what r u saying

OpenStudy (anonymous):

use this formula: sec^2x-cosec^2x=tan^2x-cot^2x

OpenStudy (anonymous):

so what's the answer?

OpenStudy (anonymous):

a^6 - b^6 = (a³)² - (b³)² = (a³+b³)(a³-b³) = (a+b)(a²-ab+b²)(a-b)(a²+ab+b²) = (a+b)(a-b)(a²+ab+b²)(a²-ab+b²) here a is sec and b is cosec.............put their values in the last

OpenStudy (anonymous):

did u got what i m trying to tell u ?

OpenStudy (anonymous):

ok, thanks for the help :)

OpenStudy (anonymous):

can i add you in Fb?

OpenStudy (anonymous):

did u got what i m trying to tell u ......if yes the click the best response if think i hav given u the best response??????/

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

thx fr awarding me d medal and becoming my fan

OpenStudy (anonymous):

it's my pleasure :)

OpenStudy (anonymous):

r u sending me req................????????

OpenStudy (anonymous):

in Fb?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

that?

OpenStudy (anonymous):

yes profile pic is mine did u gt that ?

OpenStudy (anonymous):

yes, I sent you request

OpenStudy (anonymous):

ok.